40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 1051     Entry time: Thu Oct 16 09:44:49 2008
Author: Yoichi 
Type: Update 
Category: PSL 
Subject: Bad cable for FSS 
Yesterday arount 1:30PM, we lost the LO signal for the FSS.
I found it was caused by a bad cable connecting from the peter's RF oscillator box to the LO of the FSS.
I temporarily replaced it with a BNC cable of comparable length.
ELOG V3.1.3-